OpenStudy (anonymous):

Solve using the quadratic formula. (answer in simplest exact values). 3x^2-2=-4x

OpenStudy (anonymous):

I have gotten up to x=-4± √40/6 ... just unclear on how to simplify further

OpenStudy (anonymous):

Simplify the radical. You can divid a perfect square out of it.

OpenStudy (anonymous):

in other words \[40=4\times 10\] so \[\sqrt{40}=\sqrt{4\times 10}=\sqrt4\sqrt{10}=2\sqrt{10}\]

OpenStudy (anonymous):

factor out the common factor of \(2\) in the numerator then cancel

OpenStudy (anonymous):

thank you !!

OpenStudy (anonymous):

When is Satellite wrong :\